# SAGE: Saliency-Guided Contrastive Embeddings

### v1

@@ 1.1 Background
Saliency-based training incorporates saliency maps, often derived from human experts, into the training of neural networks [4, 8, 23]. Methods focus on guiding models to salient regions of the input image, away from spurious features that can be correlated with class labels, but otherwise erroneous in solving the overarching task. Saliency maps generated by the model’s internal mechanisms are compared with the human saliency map, where the model is penalized for its deviation from the human saliency map. Incorporating such human oversight into model training is often necessary and is motivated by the severity of mission critical tasks or to challenges associated with model training. Examples of the former include high-risk domains where humans must be within the loop. The former includes open-set scenarios, highly imbalanced training data, or a limited number of training samples. Thus, saliency-based techniques are often found within the context of medicine, biometrics, robotics, and autonomous vehicles.
@@ 1.2 Motivation
These techniques, however, are constrained by their ability to convey information accurately, as this forms the underlying representation that guides the network during training. Thus, saliency maps form an upper bound on the potential for saliency-based methods. Mounting research suggests that model saliency may be unreliable, where saliency maps are not always reflective of the model’s classification mechanism [2]. If the model’s internal saliency map is compromised, any effort to steer the model towards salient features is severely undermined, or potentially outright misleading. Others have demonstrated that the means of generating model saliency (e.g. activations, gradients, etc.) can be easily manipulated, which implies that these attacks thwart the use of saliency in model training. Furthermore, saliency maps are limited in their ability to convey model outputs due to their low resolution. Their reliance on internal mechanisms constrains the saliency map’s resolution to a mere fraction of the input resolution (e.g. 3% of the input). Although this may be managed from rescaling or projections [12], it does so by improving visualization quality, not in the quantity of information portrayed. Together, these challenges cripple the validity of saliency-guided training.
@@ 2.1 Interpretable and Explainable AI
Since the conception of neural networks, researchers have sought to understand model decision-making, as a highly competent classification model is futile in high-risk tasks if human operators cannot interpret its decisions. As a result, explainable, interpretable, and trustworthy AI has spawned a copious of research methods devised to visually indicate “where the model was looking” when it arrived at a certain classification decision. Within classification models, two main approaches have been explored to extract pertinent information from the model: (a) “white-box” and (b) “black-box” approaches. “White-box” techniques utilize the model’s internal mechanisms such as its feature map activations [39, 36, 34, 35] or gradients [29, 6] to produce a model saliency map. Other methods rely on principle component analysis, image-space perturbations, or second-order gradients to formulate more fine-grained visuals. Investigations into the fidelity and trustworthiness of saliency maps have led to questions of their true reflection of model intuition, leading many researchers to either rectify these observations [13], or turn towards an entirely different paradigm of model salience altogether. “Black-box” approaches look beyond the internal mechanisms and instead observe the system as a whole. These methods opt out of utilizing the model’s internal mechanisms (e.g. activations, gradients), and instead build their saliency maps by modifying the input and measuring the differences within the output [27].
@@ 2.2 Human Saliency-guided Training
Human saliency-guided training aims to integrate human-salient information provided by saliency maps into model training, which has shown to raise model generalization [4, 8], guide the networks towards salient-features, reduce the number of training samples [8], create more human-interpretable outputs for human operators [23], and improve model convergence during training [8]. Due to their effort to align with human experts, saliency-based methods also hold several advantages within global efforts regarding AI governance and regulatory strategies [7].
Early works in integrating saliency into model training utilized class activation mappings (CAM) [4, 40]. Boyd et al. proposed CYBORG, a saliency-guided loss that generates CAMs during training penalizes the model from deviating from the human saliency map [4]. GG-CAM deviates from CYBORG in that it integrates CAMs into training using a modified classification embedding [40]. Though far less sophisticated in its saliency generation process, it has proven crucial in saliency-based methods over more complex methods due to its ability to generate model saliency in batches on a single forward pass. Other methods may require sharper visualizations by incorporating gradients, but at the computational cost of backpropagating through the network several times. Additionally, visualization can be steered off path due to their reliance on noisy and batch-dependent gradients.
Similar to CYBORG, UNET+Gaze offers a loss-based approach [19], except that model saliency is generated by a dedicated decoder rather than the classifier. Notably, the use of a dedicated decoder allows for saliency to be generated at full input resolution, unlike CYBORG [4]. These advantages have also been exploited as a pretraining strategy, where embeddings are trained from the model’s input to output, covering the entire model instead of a single layer. MENTOR decouples the training into two stages, where the model first learns to predict the saliency corresponding to each respective image, and second the model learns the classification task. MENTOR is motivated by the idea that models must first know where to look before knowing what they are looking at. Similarly, GazeGNN implements a graph neural network to impose structure on network embeddings.
@@ 3 The SAGE Approach
This section introduces SAGE, a loss-based approach to saliency-guided training through network’s embeddings that maintain outputs that follow human perception. Our method couples conventional wisdom from the existing literature in the xAI area [9]. First, the original model outputs and outputs for which the non-salient information is removed should be similar. Second, if the salient (important) regions of the input are removed, the corresponding outputs should be reflective of an image of low information (Sec. 3.1). SAGE combines these two priors into a contrastive learning approach (Sec. 3.3).
@@ 3.1 Saliency-based Augmentations
For our experiments, we use a fixed Gaussian blurring kernel size $K=(7,7)$, and a standard deviation $\sigma=10$, which controls the max intensity of the Gaussian blur. Other augmentations can be explored depending on the target domain and network architecture, such as adding random noise [26] or dimming. We then pass these images through the network to obtain model outputs (logits) and embeddings.
@@ 3.2 Saliency-based Logit Alignment
The injection of saliency-based noise (i.e. Gaussian blur) may have effects on output logit space, skewering the probability distribution in ways that may obfuscate the metric to assess both distributions. Specifically, the noise may induce changes within the scale parameters (as opposed to simple shifts, otherwise location parameters). These challenges are consistent with knowledge distillation techniques, which distill knowledge from larger capacity networks into smaller networks by similarly aligning output distributions. Consistent with these works, we use a softmax temperature scaling to increase the effect of lower probabilities, which flattens the probability distribution. Each logit is divided by $\tau$ temperature constant before the softmax is calculated, and the output is multiplied by $2^{\tau}$. Following previous work [15], we set $\tau=2$ unless stated otherwise and hold it constant throughout all experiments.
@@
We opt for JS over the more prevalent KL divergence due to its probabilistic symmetry and numerical smoothness [5, 25]. Unlike KL divergence, JS divergence is symmetric as it considers both probability distributions equally ($D(P||Q)$ and $D(Q||P)$, as opposed to simply $D(P||Q)$), which is crucial to our experiments as we are degrading the input, not reducing it to complete noise. Though the high-frequency information is removed from the input, there is still a meaningful, but degraded probability distribution that may be extracted. Second, JS is a numerically smoother compared to KL as it incorporates a mixture of both probability distributions ($M=\frac{1}{2}(P+Q)$). The inclusion of this mixture distribution is what differentiates it from Jeffrey’s divergence, which is purely a symmetric KL-divergence [17]. Given we our matching probability distributions among several loss surfaces, we find JS to serve as a built-in, soft regularization component within our loss. Preliminary experiments indicated that KL and Jeffrey’s divergence performed less favorable than JS, but still effective in increasing performance.
@@ 3.3 Saliency-Guided Contrastive Embeddings
We also guide the model towards salient features using a triplet margin contrastive loss [28]. We do so by using a forward hook at a predefined target layer in the network. For fair comparison across various architectures, we use the embeddings preceding the classification layer, which is often an adaptive pooling layer to flatten the feature maps for the classification head.
@@ 4.1 Datasets
Iris samples were samples from a larger repository of bona fide and spoof iris samples [1, 22, 14, 20, 38, 32, 21, 37, 31, 11]. We evaluate using the Iris Liveness 2020 Competition [11], consistent with prior saliency-based works [8, 10].
@@
Normal and abnormal chest X-ray radiographs were sampled from Johnson et al. [18]. Abnormal radiographs included patients diagnosed with one or several abnormalities: atelectasis (collapse of the lung), cardiomegaly (enlarged heart), pleural effusion (buildup of fluid), lung opacity (abnormal tissue density), pneumonia, edema (swelling), or support devices such as pacemakers.
@@
We follow previous works and include synthetic face detection offered by Boyd et al. [4]. Real faces are sampled from FFHQ and Celeb-HQ, whereas synthetically generated faces are provided by an assortment of StyleGAN-based generators from previous work [4].
@@ 4.2 Experimental Design
We compare our method across a wide variety of saliency-based methods. Loss-only guidance includes CYBORG [4], Unet+Gaze [19], GG-CAM [40]. MENTOR [8] is a pretraining method which guides network embeddings throughout the entire architecture using an two-stage training strategy. We also include Gaze-GNN [33], a graph neural network that incorporates saliency using embeddings from multiple encoders. Finally, we benchmark against the saliency-guided augmentations introduced in Boyd et al. [3], which includes blurring the non-salient regions of the input samples to guide models towards salient features.
For a fair comparison, each method is trained using identical configurations following previous works [4, 10, 8]. Models are initialized with ImageNet weights and trained with a batch size of $20$ for $50$ epochs using stochastic gradient descent. We set the learning rate to $0.005$ and use a weight decay of $1\times 10^{-6}$. Every 12 epochs a scheduler reduces the learning rate by a factor of $0.1$.
@@ 5.1 Generalization Performance
Interestingly, SAGE reports a more modest performance for the synthetic face detection task (0.587), followed by CYBORG (0.531), Gaze-GNN (0.524), and MENTOR (0.506). These results suggest that SAGE may not be as effective on datasets that already include high-frequency information associated with synthetic signatures, which is not appropriately captured by the saliency provided to the model and consistent with previous findings [4, 8]. Given the non-salient portions of the input may be embedded with generator-dependent signals, thereby undermining the effectiveness of the augmentation strategy and its ability to guide the network’s embeddings.
@@
We extend our analysis to include different neural network architectures to validate our initial findings (see Tab.2). SAGE raises generalization for EfficientNet-b7 backbones for Iris and Chest X-rays over existing methods [30], and DenseNet-121 for Iris [16]. Our experiments indicate that our method broadly generalizes to vision transformers, which tend to struggle in performance compared to CNNs on these datasets due to the limited supply of training data. Nevertheless, SAGE is able to increase the capacity of Swin transformer for two of the three datasets [24]. However, Swin struggled to converge for synthetic face detection, likely requiring more data or different training configurations before more analysis can be performed.
@@ 5.3 Ablation Study
We also include a small set of experiments exploring the role temperature scaling ($\tau$) may have on generalization performance (see Tab. 4). Temperature is a scaling mechanism to determine the uniformity of our logit distributions. Higher values indicate a flatter probability distribution, which relaxes the constraint on the network and encourages greater exploration of the solution space. When $\tau=1$, the distribution remains unchanged. Our experiments suggest that $\tau=2$ is optimal for both Iris and Face, whereas Chest had a slight bump in performance for a higher temperature. Additionally, it may also be an indication of the quality of information provided by the saliency maps. For example, higher temperature values may be beneficial when saliency is provided by experts (e.g. radiologists for chest X-ray anomalies), and lower values for non-experts or less experienced annotators. Interestingly, this may be a way to change the granularity of the input sample beyond the normalized image space. Our performance is on par with previous work in studying the granularity of saliency collection methods [10], which may be a fascinating connection within the hyperparameter of network training as opposed to the input space. Though not a core part of our work, we leave interesting questions related to the annotation quality for future work.

### v2

@@ 1.1 Background
Saliency-based training incorporates saliency maps, often derived from human experts, into the training of neural networks [boyd2021cyborg, crum2023mentor, linsley2018learning]. Methods focus on guiding models to salient regions of the input image, away from spurious features that can be correlated with class labels, but otherwise erroneous in solving the overarching task. Saliency maps generated by the model’s internal mechanisms are compared with the human saliency map, where the model is penalized for its deviation from the human saliency map. Incorporating such human oversight into model training is often necessary and is motivated by the severity of mission critical tasks or to challenges associated with model training. Examples of the former include high-risk domains where humans must be within the loop. The former includes open-set scenarios, highly imbalanced training data, or a limited number of training samples. Thus, saliency-based techniques are often found within the context of medicine, biometrics, robotics, and autonomous vehicles.
@@ 1.2 Motivation
These techniques, however, are constrained by their ability to convey information accurately, as this forms the underlying representation that guides the network during training. Thus, saliency maps form an upper bound on the potential for saliency-based methods. Mounting research suggests that model saliency may be unreliable, where saliency maps are not always reflective of the model’s classification mechanism [adebayo2018sanity]. If the model’s internal saliency map is compromised, any effort to steer the model towards salient features is severely undermined, or potentially outright misleading. Others have demonstrated that the means of generating model saliency (e.g. activations, gradients, etc.) can be easily manipulated, which implies that these attacks thwart the use of saliency in model training. Furthermore, saliency maps are limited in their ability to convey model outputs due to their low resolution. Their reliance on internal mechanisms constrains the saliency map’s resolution to a mere fraction of the input resolution (e.g. 3% of the input). Although this may be managed from rescaling or projections [draelos2020hirescam], it does so by improving visualization quality, not in the quantity of information portrayed. Together, these challenges cripple the validity of saliency-guided training.
@@ 2.1 Interpretable and Explainable AI
Since the conception of neural networks, researchers have sought to understand model decision-making, as a highly competent classification model is futile in high-risk tasks if human operators cannot interpret its decisions. As a result, explainable, interpretable, and trustworthy AI has spawned a copious of research methods devised to visually indicate “where the model was looking” when it arrived at a certain classification decision. Within classification models, two main approaches have been explored to extract pertinent information from the model: (a) “white-box” and (b) “black-box” approaches. “White-box” techniques utilize the model’s internal mechanisms such as its feature map activations [zhou2016learning, wang2020score, wang2019score, wang2020ss] or gradients [selvaraju2017grad, chattopadhay2018grad] to produce a model saliency map. Other methods rely on principle component analysis, image-space perturbations, or second-order gradients to formulate more fine-grained visuals. Investigations into the fidelity and trustworthiness of saliency maps have led to questions of their true reflection of model intuition, leading many researchers to either rectify these observations [fu2020axiom], or turn towards an entirely different paradigm of model salience altogether. “Black-box” approaches look beyond the internal mechanisms and instead observe the system as a whole. These methods opt out of utilizing the model’s internal mechanisms (e.g. activations, gradients), and instead build their saliency maps by modifying the input and measuring the differences within the output [petsiuk2018rise].
@@ 2.2 Human Saliency-guided Training
Human saliency-guided training aims to integrate human-salient information provided by saliency maps into model training, which has shown to raise model generalization [boyd2021cyborg, crum2023mentor], guide the networks towards salient-features, reduce the number of training samples [crum2023mentor], create more human-interpretable outputs for human operators [linsley2018learning], and improve model convergence during training [crum2023mentor]. Due to their effort to align with human experts, saliency-based methods also hold several advantages within global efforts regarding AI governance and regulatory strategies [crum2023seriously].
Early works in integrating saliency into model training utilized class activation mappings (CAM) [boyd2021cyborg, zhu2022gaze]. Boyd et al. proposed CYBORG, a saliency-guided loss that generates CAMs during training penalizes the model from deviating from the human saliency map [boyd2021cyborg]. GG-CAM deviates from CYBORG in that it integrates CAMs into training using a modified classification embedding [zhu2022gaze]. Though far less sophisticated in its saliency generation process, it has proven crucial in saliency-based methods over more complex methods due to its ability to generate model saliency in batches on a single forward pass. Other methods may require sharper visualizations by incorporating gradients, but at the computational cost of backpropagating through the network several times. Additionally, visualization can be steered off path due to their reliance on noisy and batch-dependent gradients.
Similar to CYBORG, UNET+Gaze offers a loss-based approach [karargyris2021creation], except that model saliency is generated by a dedicated decoder rather than the classifier. Notably, the use of a dedicated decoder allows for saliency to be generated at full input resolution, unlike CYBORG [boyd2021cyborg]. These advantages have also been exploited as a pretraining strategy, where embeddings are trained from the model’s input to output, covering the entire model instead of a single layer. MENTOR decouples the training into two stages, where the model first learns to predict the saliency corresponding to each respective image, and second the model learns the classification task. MENTOR is motivated by the idea that models must first know where to look before knowing what they are looking at. Similarly, GazeGNN implements a graph neural network to impose structure on network embeddings.
@@ 3 The SAGE Approach
This section introduces SAGE, a loss-based approach to saliency-guided training through network’s embeddings that maintain outputs that follow human perception. Our method couples conventional wisdom from the existing literature in the xAI area [crum2023explain]. First, the original model outputs and outputs for which the non-salient information is removed should be similar. Second, if the salient (important) regions of the input are removed, the corresponding outputs should be reflective of an image of low information (Sec. 3.1). SAGE combines these two priors into a contrastive learning approach (Sec. 3.3).
@@ 3.1 Saliency-based Augmentations
For our experiments, we use a fixed Gaussian blurring kernel size $K=(7,7)$, and a standard deviation $\sigma=10$, which controls the max intensity of the Gaussian blur. Other augmentations can be explored depending on the target domain and network architecture, such as adding random noise [parzianello2022saliency] or dimming. We then pass these images through the network to obtain model outputs (logits) and embeddings.
@@ 3.2 Saliency-based Logit Alignment
The injection of saliency-based noise (i.e. Gaussian blur) may have effects on output logit space, skewering the probability distribution in ways that may obfuscate the metric to assess both distributions. Specifically, the noise may induce changes within the scale parameters (as opposed to simple shifts, otherwise location parameters). These challenges are consistent with knowledge distillation techniques, which distill knowledge from larger capacity networks into smaller networks by similarly aligning output distributions. Consistent with these works, we use a softmax temperature scaling to increase the effect of lower probabilities, which flattens the probability distribution. Each logit is divided by $\tau$ temperature constant before the softmax is calculated, and the output is multiplied by $2^{\tau}$. Following previous work [hinton2015distilling], we set $\tau=2$ unless stated otherwise and hold it constant throughout all experiments.
@@
We opt for JS over the more prevalent KL divergence due to its probabilistic symmetry and numerical smoothness [cai2022distances, nielsen2019jensen]. Unlike KL divergence, JS divergence is symmetric as it considers both probability distributions equally ($D(P||Q)$ and $D(Q||P)$, as opposed to simply $D(P||Q)$), which is crucial to our experiments as we are degrading the input, not reducing it to complete noise. Though the high-frequency information is removed from the input, there is still a meaningful, but degraded probability distribution that may be extracted. Second, JS is a numerically smoother compared to KL as it incorporates a mixture of both probability distributions ($M=\frac{1}{2}(P+Q)$). The inclusion of this mixture distribution is what differentiates it from Jeffrey’s divergence, which is purely a symmetric KL-divergence [jeffreys1998theory]. Given we our matching probability distributions among several loss surfaces, we find JS to serve as a built-in, soft regularization component within our loss. Preliminary experiments indicated that KL and Jeffrey’s divergence performed less favorable than JS, but still effective in increasing performance.
@@ 3.3 Saliency-Guided Contrastive Embeddings
We also guide the model towards salient features using a triplet margin contrastive loss [schroff2015facenet]. We do so by using a forward hook at a predefined target layer in the network. For fair comparison across various architectures, we use the embeddings preceding the classification layer, which is often an adaptive pooling layer to flatten the feature maps for the classification head.
@@ 4.1 Datasets
Iris samples were samples from a larger repository of bona fide and spoof iris samples [casia-database, Sung_OE_2007, Galbally_ICB_2012, Kohli_ICB_2013, Yambay_ISBA_2017, Trokielewicz_IVC_2020, Kohli_BTAS_2016, Wei_ICPR_2008, Trokielewicz_BTAS_2015, das2020iris]. We evaluate using the Iris Liveness 2020 Competition [das2020iris], consistent with prior saliency-based works [crum2023mentor, crum2024grains].
@@
Normal and abnormal chest X-ray radiographs were sampled from Johnson et al. [johnson2019mimic]. Abnormal radiographs included patients diagnosed with one or several abnormalities: atelectasis (collapse of the lung), cardiomegaly (enlarged heart), pleural effusion (buildup of fluid), lung opacity (abnormal tissue density), pneumonia, edema (swelling), or support devices such as pacemakers.
@@
We follow previous works and include synthetic face detection offered by Boyd et al. [boyd2021cyborg]. Real faces are sampled from FFHQ and Celeb-HQ, whereas synthetically generated faces are provided by an assortment of StyleGAN-based generators from previous work [boyd2021cyborg].
@@ 4.2 Experimental Design
We compare our method across a wide variety of saliency-based methods. Loss-only guidance includes CYBORG [boyd2021cyborg], Unet+Gaze [karargyris2021creation], GG-CAM [zhu2022gaze]. MENTOR [crum2023mentor] is a pretraining method which guides network embeddings throughout the entire architecture using an two-stage training strategy. We also include Gaze-GNN [wang2024gazegnn], a graph neural network that incorporates saliency using embeddings from multiple encoders. Finally, we benchmark against the saliency-guided augmentations introduced in Boyd et al. [boyd2022human], which includes blurring the non-salient regions of the input samples to guide models towards salient features.
For a fair comparison, each method is trained using identical configurations following previous works [boyd2021cyborg, crum2024grains, crum2023mentor]. Models are initialized with ImageNet weights and trained with a batch size of $20$ for $50$ epochs using stochastic gradient descent. We set the learning rate to $0.005$ and use a weight decay of $1\times 10^{-6}$. Every 12 epochs a scheduler reduces the learning rate by a factor of $0.1$.
@@ 5.1 Generalization Performance
Interestingly, SAGE reports a more modest performance for the synthetic face detection task (0.587), followed by CYBORG (0.531), Gaze-GNN (0.524), and MENTOR (0.506). These results suggest that SAGE may not be as effective on datasets that already include high-frequency information associated with synthetic signatures, which is not appropriately captured by the saliency provided to the model and consistent with previous findings [boyd2021cyborg, crum2023mentor]. Given the non-salient portions of the input may be embedded with generator-dependent signals, thereby undermining the effectiveness of the augmentation strategy and its ability to guide the network’s embeddings.
@@
We extend our analysis to include different neural network architectures to validate our initial findings (see Tab.2). SAGE raises generalization for EfficientNet-b7 backbones for Iris and Chest X-rays over existing methods [tan2019efficientnet], and DenseNet-121 for Iris [huang2017densely]. Our experiments indicate that our method broadly generalizes to vision transformers, which tend to struggle in performance compared to CNNs on these datasets due to the limited supply of training data. Nevertheless, SAGE is able to increase the capacity of Swin transformer for two of the three datasets [liu2021swin]. However, Swin struggled to converge for synthetic face detection, likely requiring more data or different training configurations before more analysis can be performed.
@@ 5.3 Ablation Study
We also include a small set of experiments exploring the role temperature scaling ($\tau$) may have on generalization performance (see Tab. 4). Temperature is a scaling mechanism to determine the uniformity of our logit distributions. Higher values indicate a flatter probability distribution, which relaxes the constraint on the network and encourages greater exploration of the solution space. When $\tau=1$, the distribution remains unchanged. Our experiments suggest that $\tau=2$ is optimal for both Iris and Face, whereas Chest had a slight bump in performance for a higher temperature. Additionally, it may also be an indication of the quality of information provided by the saliency maps. For example, higher temperature values may be beneficial when saliency is provided by experts (e.g. radiologists for chest X-ray anomalies), and lower values for non-experts or less experienced annotators. Interestingly, this may be a way to change the granularity of the input sample beyond the normalized image space. Our performance is on par with previous work in studying the granularity of saliency collection methods [crum2024grains], which may be a fascinating connection within the hyperparameter of network training as opposed to the input space. Though not a core part of our work, we leave interesting questions related to the annotation quality for future work.